A sample Of ammonium phosphate, (NH 4 P04, contains 6 moles of hydrogen atoms. The
number of moles Of oxygen atoms in the sample is:
1
Expert's answer
2020-11-19T08:51:22-0500

No. of H-atoms =12

No. of O-atoms =4

No. of moles of hydrogen atoms = 6

No. of moles of oxygen atoms = nO

∴ so, ratio of no. of atoms =  ratio of no. of moles

"\frac{12}{4}" = "\frac{6}{n_o}"


nO = "\frac{24}{12}"


nO = 2